OpenStudy (anonymous):

Solve the equation. 4(x - 1.6) = 9.3

OpenStudy (chaise):

4(x - 1.6) = 9.3 We must expand the brackets, when we expand the brackets we get: 4x - 6.4 = 9.3 We must now plus 6.4 each side of the equation. 4x = 15.7 x = 15.7/4 x = 3.925

OpenStudy (anonymous):

hey sunil this side i have problem in math. i want to solve this math if x+1/x=3 then the value of (x3+1/x)/x2-x+1 is =? solve this math